OpenStudy (anonymous):

Is this proper or improper? integral 0 to 2 x^-2/5 dx

OpenStudy (anonymous):

I got 5x^3/5 over 3 + C so will it be proper or improper ?

OpenStudy (anonymous):

@Euler271

OpenStudy (anonymous):

\[\frac{ 5x ^{3/5} }{ 3} + C \]

OpenStudy (anonymous):

it is improper since the function you are integrating (regardless of if integral) is not defined at 0. this means you have to take: \[2 \lim_{t->0+} \int\limits_{t}^{2} \frac{ dx}{ x^{2/5} }\]

OpenStudy (anonymous):

ignore the 2 before the limit , lol

OpenStudy (anonymous):

Okay lol thank you!